vipi1:

Bed Day offers savings of $10, $25, $50 and $100 off on their already low price on bedroom suites. Select a coupon and save the amount on the coupon. There are a total of 100 coupons with 70, 20, 8 and 2 containing $10, $25, $50 and $100, respectively. 3. Determine the average saving.

vipi1:

Pls help me to find the standard deviation also.

itsme:

were finding the average

Hoodmemes:

To solve this I'm pretty sure u wanna multiply the coupons in order by the how much the coupon contains in order. "70, 20, 8 and 2 containing $10, $25, $50 and $100, " Lets (Multiply) (70) by (10) \[70\times10=700\] Now (Multiply) (20) by (25) \[20\times25=500\] U wanna (Multiply) (8) by (50) \[8\times50=400\] We gonna now (Multiply) (2) times (50) \[2\times100=200\] U wanna now (Add) (700) (500) (400) (200) together. I'd say u simply just wanna add (700) and (500),then (400) and (200) and add the sum of those two together. \[700+500=1200\] \[400+200=600\] \[1200+600=1800\] Now as we can see it the problem it states "There are a total of 100 coupons" Basically we wanna (Divide) (1800) by (100) \[1800\div100=18\] (18) can also be written as (18.00) So the average saving would be (18.00) dollars.